# SAGE: Spatial-visual Adaptive Graph Exploration for Visual Place Recognition

### v2

@@ A.4 Compared Methods Details
EDTFormer (jin2025edtformer)https://github.com/Tong-Jin01/EDTformer. This method introduces EDTformer, a simplified transformer decoder architecture that utilizes a set of learnable queries to efficiently decode and aggregate crucial information from image features for robust place recognition. Furthermore, it enhances the DINOv2 backbone with a novel Low-rank Parallel Adaptation (LoPA) method, which enables highly memory and parameter-efficient fine-tuning to deliver high performance with lower training costs.
@@
EffoVPR (tzachor2024effovpr). This method introduces EffoVPR, which effectively utilizes a foundation model by extracting powerful local descriptors from its internal self-attention layers for a highly effective re-ranking process, even in a zero-shot setting. Furthermore, its single-stage approach achieves state-of-the-art performance with exceptionally compact global features by simplifying training to fine-tune only the model’s final layers, thus eliminating the need for external aggregation modules.
@@ A.5 Limitations & Discussions & Future Work
Limitations and Discussions. Despite its strong performance, SAGE has some limitations. First, its performance can be compromised in highly dynamic scenes with significant, rapid occlusions (e.g., heavy traffic), as the model prioritizes static background features which may be temporarily obscured. Second, the Online Graph Construction introduces a computational step per epoch. While our experiments show this overhead is marginal on current benchmarks, the process could become time consuming for extremely large-scale datasets. This represents a trade-off between training efficiency and the adaptability of the sampling strategy. Third, the methodology’s effectiveness relies on the availability of reasonably accurate geographic coordinates during training. In scenarios with noisy or sparse GPS data, learning fine-grained spatial distinctions could be challenging.
Future Work. First, to better handle dynamic scenes, future work could integrate powerful foundation models like the Segment Anything Model (SAM) (SAM) to explicitly identify and mask transient objects. This would allow the model to focus purely on the stable, discriminative background context. Second, to enhance robustness against extreme domain shifts, we plan to incorporate multi-modal cues (han2025multimodal; xu20253d). Fusing visual features with semantic information from segmentation maps could provide complementary signals for more robust matching. Finally, the core principle of adaptive graph-based sampling holds promise beyond VPR. Applying this dynamic, “slow thinking” paradigm to other deep metric learning tasks, such as person re-identification or fine-grained image retrieval, could be a fruitful area of research.

### v3

@@ A.4 Compared Methods Details
EDTFormer (Jin et al., 2025b) [^16] https://github.com/Tong-Jin01/EDTformer. This method introduces EDTformer, a simplified transformer decoder architecture that utilizes a set of learnable queries to efficiently decode and aggregate crucial information from image features for robust place recognition. Furthermore, it enhances the DINOv2 backbone with a novel Low-rank Parallel Adaptation (LoPA) method, which enables highly memory and parameter-efficient fine-tuning to deliver high performance with lower training costs.
@@
EffoVPR (Tzachor et al., 2025). This method introduces EffoVPR, which effectively utilizes a foundation model by extracting powerful local descriptors from its internal self-attention layers for a highly effective re-ranking process, even in a zero-shot setting. Furthermore, its single-stage approach achieves state-of-the-art performance with exceptionally compact global features by simplifying training to fine-tune only the model’s final layers, thus eliminating the need for external aggregation modules.
@@ A.5 Limitations & Discussions & Future Work
Limitations and Discussions.
While achieving strong performance, we identify a few areas for further exploration. First, navigating highly dynamic scenes with rapid occlusions (e.g., as shown in Fig. 9) remains an ongoing challenge, as reliable background cues may be temporarily obscured. Second, while the per-epoch overhead of Online Graph Creation is entirely acceptable on current benchmarks, scaling to extraordinarily large datasets might incur increased computational demands, representing a conscious balance between efficiency and adaptability. Third, our graph construction benefits from reasonably accurate geographic coordinates. In scenarios with highly noisy or sparse GPS data, learning granular spatial distinctions may require further robust adaptations.
Future Work.
First, to better handle dynamic scenes, future work could integrate foundation models like SAM and CLIP to explicitly mask transient objects, encouraging the model to focus on stable background contexts. Additionally, exploring multi-view 3D representations offers a potential direction. By reasoning across multiple viewpoints and aggregating geometry-grounded features, the framework could capture richer spatial structures to enhance generalization in complex environments.
Second, to enhance robustness against extreme domain shifts (e.g., historical vs. modern images), we plan to incorporate broader multimodal cues (Ou et al., 2025; Xu et al., 2025). Fusing visual features with textual metadata, 3D contextual reasoning, or topological graph structures could provide complementary and rich semantic signals for more resilient matching across varying domains.
Finally, the core principle of adaptive graph-based sampling holds promise beyond VPR. Applying this dynamic, “slow thinking” paradigm to other deep metric learning tasks, such as person re-identification or fine-grained image retrieval, could be a fruitful area of research.